OpenStudy (erinweeks):

Suppose you are playing a game with two number cubes. Let A represent rolling 2, 3, or 4, and B represent rolling 1, 5, or 6. The probability of A is 1/2 and the probability of B is 1/2

OpenStudy (erinweeks):

a. simplify \[(\frac{ 1 }{ 2 } A + \frac{ 1 }{ 2 } B )^{2}\] b. What is the probability that one number cube shows 2, 3, or 4, and the other shows 1, 5, or 6?

OpenStudy (anonymous):

a. jus plug in the values b. when u want BOTH, the probabilites are multiplied

OpenStudy (erinweeks):

what are the values ?? Im confused !

OpenStudy (anonymous):

there are 4 possibilities first A, second A first A, second B first B, second A first B, second B each has a chance of 1/4 the second and third are the situation you're looking for. the chance of the second and third possiblity summed is 1/2

OpenStudy (anonymous):

values: A is 1/2 and B is 1/2

OpenStudy (erinweeks):

i wrote a . 1/4 , put 1/2 in for a and b and multiply 1/2 by 1/2 and b. it is a 25% chance that the cube shows 2, 3, or 4 and 25% chance that the other shows 1, 5, 6 and i got it wrong !

OpenStudy (anonymous):

a. (1/2A+1/2B)2 = (1/4+1/4)2 b. it is a 50% chance that the cube shows 2, 3, or 4 and 50% chance that the other shows 1, 5, 6 1/2 implies 50% answer: it is a 25% chance that both happen together

OpenStudy (erinweeks):

a. so thats the answer? you just multiply 1/2 * 1/2 on both sides .. b. so i kin of did it right justt didnt do the 50%?

OpenStudy (anonymous):

a. (1/2A+1/2B)2 = (1/4+1/4)2 = 1/4 b. yes